OpenStudy (anonymous):

The seventh term of the sequence -2, -6, -10, -14, ... is _____. -26 -24 -22 -20

OpenStudy (anonymous):

this is an arithmetic progression an=a1+(n-1)d

OpenStudy (anonymous):

n is 7 a1=-2 and d is the difference between two consecutive terms...

OpenStudy (anonymous):

so -20 is not the answer

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

than what will it be @ghass1978

OpenStudy (anonymous):

d=-6-(-2)=-4 just plugin the equation i gave 2...a7=a1+(7-1)(-4)...

OpenStudy (anonymous):

-26 is the answer